A force F = 1.8 i + 2.5 j N is applied at the point x = 3.0 m, y = 0 m.

(a) Find the torque about the origin.
k N
Answer #1

(a)F = 1.8 i + 2.5 j N

torque=(3 i)x(1.8 i + 2.5 j )=7.5 k N-m

(a)F = 1.8 i + 2.5 j N

torque=((3-(-1.3))i+(0-2.4))x(1.8 i + 2.5 j )=(4.3i - 2.4 j )x(1.8 i + 2.5 j ) = (4.32+ 10.75) k N-m=15.07 k N-m